Outback Outfitters sells recreational equipment. One of the company’s products, a small camp stove, sells for $140 per unit. Variable expenses are $98 per stove, and fixed expenses associated with the stove total $205,800 per month.

Required:

1. What is the break-even point in unit sales and in dollar sales?

Break-even point in unit sales
Break-even point in dollar sales   

2. If the variable expenses per stove increase as a percentage of the selling price, will it result in a higher or a lower break-even point? (Assume that the fixed expenses remain unchanged.)

3. At present, the company is selling 12,000 stoves per month. The sales manager is convinced that a 10% reduction in the selling price would result in a 25% increase in monthly sales of stoves. Prepare two contribution format income statements, one under present operating conditions, and one as operations would appear after the proposed changes.

Outback Outfitters
Contribution Income Statement Present Proposed
12,000 Stoves Stoves
Total Per unit Total Per unit
0 $0 0 $0
$0 $0

4. Refer to the data in Required 3. How many stoves would have to be sold at the new selling price to attain a target profit of $73,000
Answer #1

Selling price per unit = $140

Variable cost per unit = $98

Fixed expenses = $205,800

Contribution margin per unit = Selling price per unit - Variable cost per unit

= 140-98

= $42

Contribution margin ratio = Contribution margin per unit/ Selling price per unit

= 42/140

= 30%

1.

Break even point ( in units) = Fixed costs / Contribution margin per unit

= 205,800/42

= 4,900 units

Break even point ( in dollar) = Fixed costs / Contribution margin ratio

= 205,800/30%

= $686,000

Break-even point in unit sales 4,900 units
Break-even point in dollar sales $686,000

2.

If variable expense per stove increase as a percentage of selling price, it will result in higher Break even point.

3.

Present quantity sold = 12,000 stoves

Reduction in selling price = 10%

= 140 x 10%

= $14

Proposed selling price = 140-14

= $126

Increase in sales = 25%

= 12,000 x 25%

= 3,000 units

Proposed selling quantity = 12,000+3,000

= 15,000 units

Outback Outfitters
Contribution Income Statement Present Proposed
12,000 Stoves 15,000 Stoves
Total Per unit Total Per unit
Sales (12,000 x 140)=1,680,000 140 (15,000 x 126) =1,890,000 126
Variable cost -(12,000 x 98)=1,176,000 -98 -(15,000 x 98)=1,470,000 -98
Contribution margin 504,000 $42 420,000 $28
Fixed cost -205,800 -205,800
Operating Income $298,200 $214,200

4.

Number of units to be sold for target profit = ( Fixed cost + Target profit)/Contribution margin per unit

= (205,800+73,000)/42

= 6,638 units

Kindly comment if you need further assistance.

Thanks‼!